# DeepSpeak Dataset v1.0

### v2

@@ Abstract
We describe a large-scale dataset–DeepSpeak–of real and deepfake footage of people talking and gesturing in front of their webcams. The real videos in this first version of the dataset consist of $1717$ hours of footage from $220220$ diverse individuals. Constituting more than $2626$ hours of footage, the fake videos consist of a range of different state-of-the-art face-swap and lip-sync deepfakes with natural and AI-generated voices. We expect to release future versions of this dataset with different and updated deepfake technologies. This dataset is made freely available for research and non-commercial uses; requests for commercial use will be considered.
@@ 1 Introduction
Large-scale and diverse datasets are critical to the academic literature to aid in both system development and evaluation. We describe an audio and video dataset designed to aid the digital-forensic community. This dataset consists of $4444$ hours of real and deepfake video of people talking and gesturing. The real videos were recorded with consent from the participants. The deepfakes consists of face-swap deepfakes (with three variants), lip-sync deepfakes (from two different synthesis engines), and audio deepfakes spliced into a subset of the lip-sync deepfakes. Shown in Table 1 is a comparison of the characteristics of our dataset with recent (past six years) datasets.
We describe the details of the collecting of the real video and how the various deepfake audio and video are created. The full dataset is available by request at https://huggingface.co/datasets/faridlab/deepspeak_v1.
@@
FaceForensics  [16] 2018 ? 1,004 N $\checkmark$ - - - 2,008
FaceForensics++ [17] 2019 ? 1,000 N $\checkmark$ - $\checkmark$ - 4,000
@@
DeepSpeak v1.0 2024 220 6,226 Y $\checkmark$ $\checkmark$ - $\checkmark$ 6,799
Table 1: A comparison of forensic-themed public datasets. Although not the most informative metric, we report original and deepfake footage as number of videos for consistency with previous published datasets.
@@ 2 Data Collection
Recordings of 220 participants were crowd-sourced through the Prolific research recruitment platform. Participants were asked to give their consent for including their recordings, without any other identifying information, in a public dataset [^1] The precise consent language was: “This dataset will be used for research purposes for detecting deepfakes. Please note that your recordings will be made public in a dataset, but no other identifying information will be shared outside of our research group.”. Participants were selected from a stratified sample ensuring equal distribution of gender, with all participants being native English speakers and U.S. residents. The demographics of the participants were as follows (some participants identified with more than one race/ethnicity):
@@
Each participant completed a survey in which they were instructed to record themselves responding to 32 separate prompts (see Appendix A). Participants were paid \$7 for their time. The first two prompts were used only for voice-clone training data, and were omitted from the final dataset (see Section 3). The remaining prompts were divided into four categories: (1) standardized scripted responses in which each participant read the same prompt; (2) randomized scripted responses in which participants read a randomized prompt; (3) unscripted responses in which participants responded to questions; and (4) actions in which participants performed simple actions.
Participants then responded to four open-ended unscripted questions and asked to aim for a response that was close to 30 seconds in length. Participants then responded to give quick-fire questions in which they repeated the question and provided a short response. See Appendix A for all these prompts.
@@
Following the scripted and unscripted responses, participants were asked to perform four simple visual actions (1) wave hand in front of face while counting 1,2,3; (2) look down and right, straight down, and down and left, each time holding and counting 1,2,3; (3) look up and right, straight up, and up and left, each time holding and counting 1,2,3; (4) lean towards the camera without counting.
We manually removed from the final dataset any invalid responses. This included ensuring that only a single person was visible in the video, the scene was reasonably well lit, and each submitted video clip contained a valid audio and visual track. The audio/video recordings were converted from their initial .webm format to .mp4 and organized by prompt type. Videos were encoded using the H.264 codec with the libx264 library. Audio-only .wav tracks were also extracted from each response for voice synthesis (see Section 3).
@@ 3 Audio Generation
Two scripted prompts of average length of 30 seconds each were used as inputs for voice cloning using ElevenLabs’ voice cloning API. Participants were first asked to record themselves reading 10 consecutive phonetically-rich sentences, sourced from List 1 of the standard Harvard Sentences [18]. Participants were then asked to repeat the standard elicitation paragraph from the Speech Accent Archive, a phonetically comprehensive passage comprising a breadth of vowels and consonants [20]. These responses were used for the purpose of voice cloning and are omitted from the final video dataset.
Using each participant’s cloned voice, a synthetic audio was created in their voice saying the same thing as in the original audio/video. For the unscripted responses, the original audio was transcribed using OpenAI’s Whisper, and for the scripted responses, we assumed that the participant correctly read the script. These text transcriptions were then provided to ElevenLabs’ API to generate matching synthetic voices.
@@ 4 Video Generation
[…]
We next describe the different models/configurations for the creation of face-swap and lip-sync deepfakes.
@@ 4.1 Face-Swap
Face-swap deepfakes are created by replacing – eyebrow to chin and cheek to cheek – the original identity in a video with a new identity. In order to create a visually compelling face-swap deepfake, each identity in the full dataset is paired with a perceptually similar face as follows. Each identity is first represented by a mean CLIP embedding [14] of five randomly extracted video frames. For each identity a unique matched identity is assigned using the agglomerative clustering algorithm with cosine distance and cluster size constraint from the scikit-learn library [^2] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.AgglomerativeClustering.html. Representative examples of these pairs are shown in Figure 2.
@@
Figure 2: Four representative examples of matched identities.
FaceFusion: The first face-swap configuration invokes the FaceFusion [19] library with its default parameters. FaceFusion, built on InsightFace [15], localizes the face in the original video, performs 3D landmark estimation of the face, and superimposes the appropriately transformed matched face. This is followed by a set of post-processing steps to hide edge artifacts and improve temporal consistency.
@@
FaceFusion Live: The third face-swap configuration wraps the same FaceFusion configuration in a simulated live input streaming environment. Because our hardware (a Tesla T4 GPU) cannot generate video frames in real-time ($2424$ to $3030$ frames per second (fps)), the effective frame-rate of the generated video is decreased to approximately $1212$ fps
@@ 4.2 Lip-Sync
Three types of lip-sync deepfakes were created using two different models. Given an original video and associated audio, the first lip-sync deepfake consists of an audio of the same identity from a different video (i.e., the audio and video are now mismatched). The second lip-sync consists of an AI-generated voice of the same identity (Section 3) with a transcript taken from a different video. The third lip-sync consists of an audio from the matched identity (see above).
@@ 5 Discussion
Acknowledgements
This work was partially supported by Google/YouTube and the University of California Noyce Initiative. We are grateful to ElevenLabs (https://elevenlabs.io) for API access for large-scale voice generation.
@@ A.2 Standardized Scripted Prompts
Table 3: Standardized scripted prompts, consisting of ten separate sentences.
@@ A.3 Voice Cloning Prompts
Table 4: Voice cloning input prompts, consisting of ten consecutive sentences and one continuous paragraph.
@@ A.4 Randomized Scripted Prompts
Table 5: 728 randomized scripted prompts, for which each participant only responded to 10.

### v3

# The DeepSpeak Dataset

@@ Abstract
We describe a large-scale dataset – DeepSpeak – of real and deepfake footage of people talking and gesturing in front of their webcams. The real videos in this dataset consist of a total of 50 hours of footage from 500 diverse individuals. Constituting more than 50 hours of footage, the fake videos consist of a range of different state-of-the-art avatar, face-swap, and lip-sync deepfakes with natural and AI-generated voices. We are regularly releasing updated versions of this dataset with the latest deepfake technologies. This preprint describes the construction of versions 1.0, 1.1, and 2.0. This dataset is made freely available for research and non-commercial uses; requests for commercial use will be considered.
@@ 1 Introduction
Large-scale and diverse datasets are critical to the academic literature to aid in both system development and evaluation. We describe an audio and video dataset designed to aid the digital-forensic community [^1] We named our dataset DeepSpeak before the release of the DeepSeek AI model. This dataset consists of 100 hours of real and deepfake video of people talking and gesturing. The real videos were recorded with consent from the participants. The deepfakes consist of avatar (from three generators), face-swap deepfakes (with multiple variants from three generators), lip-sync deepfakes (from four generators), and audio deepfakes (from three generators) spliced into a subset of the lip-sync deepfakes. Shown in Table 1 is a comparison of the characteristics of the different versions of our dataset as compared to recent (past seven years) datasets.
We describe the details of the collection of the real video and how the various deepfake audio and video are created. Each version of the dataset is available by request at the following locations:
@@
FaceForensics  [25] 2018 NA 1,004 N $\checkmark$ - - - 2,008
FaceForensics++ [26] 2019 NA 1,000 N $\checkmark$ - $\checkmark$ - 4,000
@@
DF40 [35] 2024 NA NA N - 100,000+
DeepSpeak v1.0 2024 219 6,226 Y $\checkmark$ $\checkmark$ - $\checkmark$ 6,799
DeepSpeak v1.1 2025 220 6,667 Y $\checkmark$ $\checkmark$ - $\checkmark$ 6,796
DeepSpeak v2.0 2025 280 9,376 Y $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 7,209
Table 1: A comparison of forensic-themed public datasets. Although not the most informative metric, we report original and deepfake footage as number of videos for consistency with previous published datasets (NA: not available).
@@ 2 Data Collection
Participants were crowd-sourced through the Prolific research recruitment platform. Participants were asked to give their consent for including their recordings, without any other identifying information, in a public dataset [^2] The precise consent language was: “This dataset will be used for research purposes for detecting deepfakes. Please note that your recordings will be made public in a dataset, but no other identifying information will be shared outside of our research group.”. Participants were selected from a stratified sample ensuring equal distribution of gender, with all participants being native English speakers and U.S. residents.
Each participant completed a survey in which they were instructed to record themselves responding to 32 (version 1.0) and 35 (version 2.0) separate prompts (see Appendix A). Participants were paid \$7 for their time. The first two prompts were used for voice-clone training data (see Section 3). The remaining prompts were divided into four categories: (1) standardized scripted responses in which each participant read the same prompt; (2) randomized scripted responses in which participants read a randomized prompt; (3) unscripted responses in which participants responded to questions; and (4) actions in which participants performed simple actions.
Participants then responded to four open-ended unscripted questions and asked to aim for a response that was close to 30 seconds in length. These were followed by quick-fire questions in which they repeated the question and provided a short response. Following the scripted and unscripted responses, participants were asked to perform simple actions using head and hand gestures. See Appendix A for the full list of prompts and scripts used.
We manually removed any invalid responses from the final dataset. This included ensuring that only a single person was visible in the video, the scene was reasonably well lit, and each submitted video clip contained a valid audio and visual track.
2.1 Version 1.0
We collected data from 220 participants for version 1.1 (version 1.0 had only 219 participants due to a corrupt file). The demographics of the participants are as follows (some participants identified with more than one race/ethnicity):
@@
Participants were asked to perform four simple visual actions: (1) wave their hand in front of their face while counting 1,2,3; (2) look down and right, straight down, and down and left, each time holding and counting 1,2,3; (3) look up and right, straight up, and up and left, each time holding and counting 1,2,3; (4) lean towards the camera without counting.
The audio/video recordings were converted from their initial .webm format to .mp4 and organized by prompt type using the H.264 codec with the libx264 library (this final video conversion is not performed in version 2.0 in order to minimize re-encoding).
Because version 1.1 corrects errors from version 1.0, we recommend using this version.
@@
2.2 Version 2.0
We collected data from 280 participants for version 2.0. The demographics of the participants were as follows (some participants identified with more than one race/ethnicity):
  • Age: Range = 18-73 years, Mean = 38 years; standard deviation = 11.6 years
  • Gender: 146 male, 129 female, 3 non-binary, 2 not provided
  • Race/Ethnicity: 203 White/Caucasian, 49 Black/African American, 20 Asian, 11 American Indian/Alaska Native, 1 Native Hawaiian/Other Pacific Islander, 10 other, 1 prefer not to say.
Participants were asked to perform the same four simple visual actions as in version 1.0: (1) wave their hand in front of their face while counting 1,2,3; (2) look down and right, straight down, and down and left, each time holding and counting 1,2,3; (3) look up and right, straight up, and up and left, each time holding and counting 1,2,3; (4) lean towards the camera without counting). An additional set of three new actions were added to this version: (5) pretending to yawn; (6) pretending to laugh for between 2 to 5 seconds; (7) clapping loudly three times, pausing for about 5 seconds between claps.
While audio/video recordings for version 1.0 were converted from their initial .webm format to .mp4 using the H.264 codec with the libx264 library, an FFmpeg “copy” command was used in version 2.0 without re-encoding to preserve recording quality.
@@ 3 Audio Generation
Participants were first asked to record themselves reading 10 consecutive phonetically-rich sentences, sourced from List 1 of the standard Harvard Sentences [27]. Participants were then asked to repeat the standard elicitation paragraph from the Speech Accent Archive, a phonetically comprehensive passage comprising a breadth of vowels and consonants [33]. These two scripted responses were used for the purpose of voice cloning, and had an average length of 30 seconds.
Using each participant’s cloned voice, a synthetic audio was created in their voice saying the same thing as in the original audio/video. For the unscripted responses, the original audio was transcribed using OpenAI’s Whisper, and for the scripted responses, we assumed that the participant correctly read the script. These text transcriptions were then provided to each voice cloning generators’ API to generate matching synthetic voices. The following generators were used for each version of the DeepSpeak dataset.
3.1 Version 1.0
Voice clones were generated using the Elevenlabs Create Voice endpoint [^3] https://elevenlabs.io/docs/api-reference/voices/add. The Create Speech API [^4] https://elevenlabs.io/docs/api-reference/text-to-speech/convert was used for speech synthesis. The eleven_multilingual_v2 model was used, with audio output returned in MP3 format at 44.1 kHz. The API was accessed in April 2024.
3.2 Version 2.0
As in version 1.0, voice clones in version 2.0 used the same two scripted prompts for all voice-cloning generators.
The ElevenLabs API was accessed as described in the previous section.
Voice clones were generated using the PlayAI Instant Voice Cloning endpoint [^5] https://docs.play.ht/reference/api-create-instant-voice-clone. The pyht Python package, with TTSOptions set to default parameters was used for speech synthesis, with audio output provided in MP3 format, at a sampling rate of 24 kHz.
Voice clones were generated using the Speechify v1 Voices API [^6] https://docs.sws.speechify.com/v1/api-reference/api-reference/tts/voices/create. The Speech API endpoint [^7] https://docs.sws.speechify.com/v1/api-reference/api-reference/tts/audio/speech was used for speech synthesis. The simba-english model was used, with output audio returned in WAV format at 48 kHz.
All API calls were made in November 2024.
@@ 4 Video Generation
We generated three types of video deepfakes: face swap, lip sync, and avatar, each of which is described next. Between version 1.0 and 2.0 of DeepSpeak, deepfake generators improved considerably. Because we are primarily interested in creating a dataset using state-of-the-art generators, we do not, in version 2.0, re-use the same generators as in version 1.0.
All source videos were re-encoded using FFmpeg with the H.264 codec at a constant rate factor of 18. The encoding present was set to slow, and the original audio stream was preserved without re-encoding.
@@ 4.1 Face-Swap
Face-swap deepfakes are created by replacing – eyebrow to chin and cheek to cheek – the original identity in a video with a new identity. In order to create a visually compelling face-swap deepfake, each identity in the full dataset is paired with a perceptually similar face as follows. Each identity is first represented by a mean CLIP embedding [22] of five randomly extracted video frames. For each identity a unique matched identity is assigned using the agglomerative clustering algorithm with cosine distance and cluster size constraint from the scikit-learn library [^8] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.AgglomerativeClustering.html. Representative examples of these pairs are shown in Figures 2 and 3.
@@
Figure 2: Four representative examples of matched identities from version 1.0.
4.1.1 Version 1.0
We used three variants of the open-source library FaceFusion as follows:
  • FaceFusion The first face-swap configuration invokes the FaceFusion [28] library with its default parameters. FaceFusion, built on InsightFace [24], localizes the face in the original video, performs 3D landmark estimation of the face, and superimposes the appropriately transformed matched face. This is followed by a set of post-processing steps to hide edge artifacts and improve temporal consistency.
@@
  • FaceFusion Live: The third face-swap configuration wraps the same FaceFusion configuration in a simulated live input streaming environment. Because our hardware (a Tesla T4 GPU) cannot generate video frames in real-time ($2424$ to $3030$ frames per second (fps)), the effective frame-rate of the generated video is decreased to approximately $1212$ fps.
4.1.2 Version 2.0
Between the construction of version 1.0 and version 2.0, FaceFusion expanded to include a larger suite of face-swap generators. For version 2.0, we used four variants from the FaceFusion library:
  • INSwapper: The first face-swap configuration invokes the INSwapper-128 model [30] with default parameters. This model leverages InsightFace [24] to localize 3D facial landmarks and superimpose a transformed face of the matched identity onto the source identity video. This version is most similar to FaceFusion in version 1.0.
  • INSwapper + CodeFormer: The second face-swap configuration extends the above INSwapper configuration by enhancing each frame of the generated deepfake using CodeFormer [40]. This model was trained to fix any misshapen teeth or lips and improve the overall photorealism of the generated face. This version is most similar to FaceFusion+GAN in version 1.0.
  • SimSwap: The third face-swap configuration invokes the SimSwap-256 model [3] with default parameters. Unlike INSwapper, which directly modifies the target face based on detected facial landmarks, SimSwap extracts a latent identity representation to guide the generation of new frames.
  • SimSwap + RestoreFormer: The fourth face-swap configuration extends the above SimSwap configuration by enhancing each frame of the generated deepfake using RestoreFormer++ [32]. This model, trained on degraded photographs, corrects any structural deficiencies and enhances the photorealism of the generated frames.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Four representative examples of matched identities from version 2.0.
@@ 4.2 Lip-Sync
Given an original video and associated audio, we create two types of lip-sync deepfakes: (1) a lip-sync deepfake with an audio of the same identity extracted from a different video (i.e., the audio and video are now mismatched); and (2) a lip-sync deepfake with an AI-generated voice of the same identity (Section 3) with a transcript taken from a different video.
4.2.1 Version 1.0
Two types of lip-sync deepfakes were created using the following models:
@@
4.2.2 Version 2.0
Two types of lip-sync deepfakes were created using the following models:
  • Diff2Lip: The first lip-sync configuration invokes the Diff2Lip model [19] with default parameters. After localizing the face in the video, Diff2Lip adds noise over the mouth region and proceeds as an audio-conditioned diffusion model. It was trained on the Voxceleb2 [5] and LRW [36] datasets.
  • LatentSync: The second lip-sync configuration invokes the LatentSync model [14] with default parameters. While LatentSync is, at its core, a diffusion model similar to Diff2Lip, it differs in two aspects. First, it encodes the audio constraint as Whisper embeddings [23] instead of chunks of audio signal. Second, it uses a noise mask delineated by facial landmarks to exactly match the shape of the face as opposed to a rectangular bounding box. LatentSync was trained on the VoxCeleb2 [5] and HDTF [38] datasets.
4.3 Avatar
Avatar deepfakes animate the head and lip movements of a static image to match a target video or audio track. Unlike face-swap and lip-sync deepfakes, which modify an existing video, avatar deepfakes generate movement from a single static image. Version 1.0 contains no avatar deepfakes because these models were still in their infancy.
4.3.1 Version 2.0
Three types of avatar deepfakes were created using the following models:
  • LivePortrait: The first avatar configuration invokes the LivePortrait model [9] with its default parameters. LivePortrait is a multi-stage model that first extracts the identity and motion embeddings, which are warped into a joint embedding that is later decoded into pixel space. This model was trained on the VoxCeleb [20], MEAD [31], RAVDESS [17], and AAHQ [16] datasets.
  • HelloMeme: The second avatar configuration invokes the HelloMeme [37] model with its default parameters. HelloMeme consists of three modules: a reference module, which extracts an identity embedding; a control module, which extracts information about the head and mouth shape; and a diffusion module, which generates the resulting video frames. This model was trained on the CelebV-HQ [34] and VFHQ [34] datasets.
  • Memo: The third avatar configuration invokes the Memo model [39] with its default parameters. Memo is a diffusion model that combines identity, voice, and emotion embeddings as diffusion constraints. It was trained on a compilation of the HDTF [38], VFHQ [34], CelebV-HQ [42], MultiTalk [29], and MEAD [31] datasets.
LivePortrait and HelloMeme take as input a single image of a person to animate with a video (and associated audio) that drives this animation. For these two generators, the avatar deepfakes contain only real audio from the original driving video. Memo takes as input a single image of a person to animate with only an audio that drives this animation. In this case, the audio can be either real or fake.
4.4 Summary
Shown in Table 2 is a breakdown of the total size in file size (GB), file counts (N), and video length (hrs). Also shown is the breakdown for a 80/20 training/testing split. Across these splits, there is no overlap in facial or voice identities.
Total Train Test
ver size (GB) size (N) size (hrs) real (N [hrs]) fake (N [hrs]) real (N [hrs]) fake (N [hrs])
1.0 40 13,025 44.3 4,902 [13.9] 5,300 [21.0] 1,324 [3.7] 1,499 [5.8]
1.1 46 13,463 48.0 5,251 [16.8] 5,299 [21.0] 1,416 [4.4] 1,497 [5.8]
2.0 124 16,585 52.7 7,513 [23.6] 5,793 [18.6] 1,863 [5.8] 1,416 [4.6]
Table 2: A breakdown of the total size (gigabytes (GB), number of files (N), and length in hours (hrs)) of each version of the DeepSpeak dataset.
@@ 5 Discussion
Acknowledgments
This work was partially supported by Google/YouTube and the University of California Noyce Initiative. We are grateful to ElevenLabs (https://elevenlabs.io) and PlayAI (https://play.ht/) for granting us API access for voice generation.
@@ 5. If I could time travel, where in the past or future would I like to go?
Version 2.0 only
Pretend to yawn, perhaps covering your mouth with your hand.
Pretend to laugh for between 2 to 5 seconds.
Clap loudly three times, pausing for about 5 seconds between claps. Please make sure your clap makes a loud sound.
@@ A.2 Standardized Scripted Prompts
Table 4: Standardized scripted prompts, consisting of ten separate sentences for version 1.0 and 2.0.
@@ A.3 Voice Cloning Prompts
Table 5: Voice cloning input prompts for version 1.0 and 2.0, consisting of ten consecutive sentences and one continuous paragraph.
@@ A.4 Randomized Scripted Prompts
Table 6: 728 randomized scripted prompts for version 1.0 and 2.0, from which each participant only responded to 10.